### 24 Feb 2024 v2

# user interface, color vision, color blindness, CVD, accessibility

Abstract
Over 300 million people who live with color vision deficiency (CVD) have a decreased ability to distinguish between colors, limiting their ability to interact with websites and software packages. User interface designers have taken various approaches to tackle the issue with most offering a high contrast mode. The Web Content Accessibility Guidelines (WCAG) outline some best practices for maintaining accessibility that have been adopted and recommended by several governments; however, it is currently uncertain how this impacts perceived user functionality and if this could result in a reduced aesthetic look. In the absence of subjective data, we aim to investigate how a CVD observer might rate the functionality and aesthetics of existing UIs. However, the design of a comparative study of CVD vs. non-CVD populations is inherently hard, therefore we build on the successful field of physiologically-based CVD models, and propose a novel simulation-based experimental protocol, where non-CVD observers rate the relative aesthetics and functionality of screenshots of 20 popular websites as seen in full color vs. with simulated CVD. Our results show that relative aesthetics and functionality correlate positively and that an operating-system-wide high contrast mode can reduce both aesthetics and functionality. While our results are only valid in the context of simulated CVD screenshots, the approach has the benefit of being easily deployable, and can help to spot a number of common pitfalls in production. Finally, we propose a AAA–A classification of the interfaces we analyzed.
keywords:
user interface, color vision, color blindness, CVD, accessibility
[…]
1 Introduction
Color vision deficiency (CVD, color blindness) is the failure or decreased ability to distinguish between colors under normal illumination. There are over 300 million people with CVD, including approx. 1 in 12 men (8%) and 1 in 250 women (0.5%) birch2012worldwide; rigden1999eye; al2001prevalence. CVD is an X-linked genetic disorder impacting both eyes with varying degrees of prevalence in different populations fakorede2022prevalence. It affects an individual’s ability to perform tasks in both personal and professional settings tagarelli2004colour.
Color is an important asset in user interface (UI) design wcag21, and while the exact impact of color is known to vary between demographics swasty2017does, applications still often rely on established conventions, such as green and red indicating ‘yes‘ and ‘no’ respectively. Objects of the same color satisfy the Gestalt principle of similarity, whereas different colors can help an object stand out or mark figure-ground articulation todorovic2008gestalt. With the ever-increasing color gamut of novel displays wen2021color, new domains are opening up in the use of color; however, some of these domains simply cannot be seen by someone with CVD.
Accessibility is the concept of making UIs equally usable by all types of users, enabling user interactions without barriers. UI designers have the option to support accessibility for CVD users pre-publication or post-publication jefferson2006accommodating, with pre-publication normally resorting to a limited and fixed color palette rigden1999eye, and post-production relying on automatic recoloring iqbal2018adaptive also known as daltonization simon2016evaluating. A hybrid, low-effort approach is to provide support for the operating system’s high contrast mode. The Web Content Accessibility Guidelines (WCAG) wcag21 outline some best practices for accessibility; however, these often only target core functionality. Another consideration for UI is perceived aesthetics. Many designers see aesthetics as inversely proportional to functionality Karvonen_2000, while other evidence points towards a positive correlation between functionality and aesthetics Ahmed_Al; kurosu1995apparent; tractinsky1997aesthetics. In the domain of CVD, a high contrast theme is rarely a top-priority feature, which could imply that people with CVD have a substantially reduced aesthetic experience. However, there is insufficient data to understand CVD users’ perceived functionality and aesthetics of UIs.
A comparative study of UI functionality and aesthetics is inherently challenging. Individuals with CVD cannot judge if a reduced-dimensionality UI they see is equally usable or aesthetic compared to a UI they could never see. Therefore, we instead built on the successful field of physiologically-based CVD simulations and asked 19 non-CVD participants to compare reference UIs to how they might appear to a CVD observer for 20 popular UIs (1449 data points in total). Specifically, we measured mean-opinion scores for functionality and the probability of maintained aesthetics. One clear limitation of our study is non-CVD users’ potential bias towards familiar full-color references, therefore our results cannot be used to conclude whether any of the UIs are fully accessible; however, comparing scores across different applications and accessibility techniques, we can investigate common pitfalls of UI design. Our main contributions can be summarized as follows:
  • Collection of subjective user data from non-CVD observers on the loss of functionality and aesthetics when seen through CVD simulation
  • Analysis of results suggesting a positive correlation between functionality and aesthetics scores, and evidence that OS-enabled high contrast mode might be detrimental
  • Publication of the dataset and participant responses
The rest of the paper is structured as follows: in Section 2, we discuss the relevant background literature, in Section 3 we describe the proposed experimental methodology, before discussing our findings in Sections LABEL:sec:res and LABEL:sec:discussion.
2 Background
2.1 CVD
Human color vision is widely considered trichromatic due to the three types of retinal cone cells (L, M, and S) that respond differently to wavelengths of light (Figure 1) wandell1995foundations. According to opponent color theory hurvich1957opponent, higher-order visual functions rely on a differential response of cones, with luminance (L+M) more prominent than chrominance channels (L-M, L+M-S). CVD can occur as a result of cone cells being absent, not working, or having an abnormal response. Severe CVD occurs when two or three types of cone cells are absent (monochromacy, achromatopsia). Less severe CVD occurs when one type of cone is absent (L: protanopia, M: deuteranopia, S: tritanopia), or all three are present, but one cell has an abnormal response (L: protanomaly, M: deuteranomaly, S: tritanomaly) huang2008enhancing. Red-green CVD, specifically protanomaly and deuteranomaly, are the most frequent marey2015ishihara, which can be intuitively explained by the spectral similarity of the M and L cones. This affects approx. 8% and 0.1–0.3% of the male and female population respectively birch2012worldwide.
Color appearance is context-dependent, highly subjective fairchild2013color, and especially as color plays such a key role in visual perception, most find it challenging to imagine how another individual (especially with CVD) might perceive the world around them. However, the physiological mechanisms of early vision are relatively well-understood, and existing models have been shown to be capable of faithfully simulating what a CVD observer might see Machado2009. We build on these models in our experimental methodology.
Figure 1: Normalised spectral sensitivity of the three types of retinal cone cells as a function of light wavelength. Note the similarity of the $M$ and $L$ cones, where minor individual differences can result in an unresolvable overlap, causing observers to have a reduced ability to differentiate between lights in the red-green region. Plotted based on Stockman and Sharpe stockman2000spectral.
2.2 WCAG
Accessibility is the concept of making UIs equally usable by all types of users, enabling all user interactions without barriers. In our work, we consider functionality to be synonymous with usability, therefore accessibility also aims to maintain functionality. In the past few decades, accessibility has become a legal requirement in several countries easton2012revisiting, with most government agencies recommending the adoption of the World Wide Web Consortium’s (W3C) Web Content Accessibility Guidelines (WCAG) easton2012revisiting; wcag21. In fact, WCAG is consulted by website and UI developers every day to enforce best practices to maintain a level playing field for all users, regardless of age, or physical or mental disabilities. WCAG has also been utilized in academic research to assess website accessibility Al-Khalifa_2010; kuzma2010accessibility; paul2023accessibility.
WCAG 2.1 wcag21, the latest framework, specifies a set of success criteria in an AAAA system, A being the least and AAA being the most restrictive. Minimum accessibility compliance is often set at AA easton2012revisiting. CVD support is addressed across several criteria, e.g.:
  • 1.4.1 Use of Color (A): encourages providing information conveyed via color through other visual means
  • 1.4.5 Images of Text (AA): encourages text being used to convey information rather than images of text
  • 1.4.11 Non-text Contrast (AA): visual presentation of UI components and graphical objects having a contrast ratio of at least 3:1
  • 1.4.6 Contrast (Enhanced) (AAA): visual presentation of text and images of text has a contrast ratio of at least 7:1.
It remains uncertain to what extent current UIs implement these recommendations. Also, while the use of contrast ratio thresholds gives designers an intuitive scheme, contrast is defined as a color difference perceived by trichromatic (non-CVD) observers, which is not trivially convertible to contrast for CVD users. Furthermore, WCAG contrast is typically calculated for background vs. foreground, which can fail to capture some impacts of CVD, such as reduced differences across UI components (Figure 2). We accept the importance and impact of WCAG, but we argue that a user study is more appropriate for CVD.
Figure 2: Illustration of a popular quiz platform’s UI processed with simulated tritanomaly Machado2009. While text vs. background contrast remains mostly unchanged, $\Delta E_{2000}$ color difference luo2001development between buttons 1 and 2 is reduced from $25.1$ to $14.1$. Similarly, $\Delta E_{2000}$ between buttons 3 and 4 is reduced from $43.0$ to $18.3$. The colors remain distinguishable, but the difference between them is less noticeable. WCAG analysis would flag an insufficient text vs. background contrast on button 3 (<3), but current guidelines do not address the reduced color difference between buttons.
2.3 Reducing the impact of CVD
To reduce the impact of CVD, WCAG encourages UI designers to convey information via alternate visual channels. This also agrees with recent research on the use of patterns sajadi2012using, and textures wu2019composition to improve accessibility.
Another widely adopted approach limits the color palette with an increased contrast between key colors rigden1999eye; jefferson2006accommodating. E.g. Windows users can select from multiple operating-system wide high contrast themes that accommodate different types of vision deficiencies. While almost all applications and websites respect the high contrast theme of the operating system, the resulting UI often looks substantially different, with seemingly reduced information for a non-CVD observer (see Figure 3). In our study, we investigate whether this strategy is viable by measuring functionality and aesthetics with the operating system’s high contrast mode enabled and disabled.
An alternative technique is daltonization, an umbrella term for image-based recoloring techniques that enhance displayed images for CVD observers iqbal2018adaptive; ribeiro2019recoloring; machado2010real. Such algorithms aim to alter colors that are isochromatic to CVD observers but are perceived differently by a non-CVD observer. Daltonization could substantially improve CVD accessibility; however, current algorithms are still limited by e.g. run-time cost, lack of temporal coherence, or the ability to accommodate different types and levels of CVD ebelin2023luminance.
2.4 Aesthetics
While accessibility and overall functionality are key features of a good UI, designers, and human-computer interaction (HCI) researchers are increasingly discovering the impact of attractiveness and beauty. An attractive UI can help make a connection between a user and the application turkyilmaz2015user. Interdisciplinary research has investigated the relevance of aesthetics, a classical philosophical field focusing on the “perception of the beautiful in nature and art” oxford_english_dictionary_notitle_2023. A particular focus has been the difference between classical and expressive aesthetics in HCI, with classical aesthetics referring to the traditional notions of an orderly and clear design (e.g. sizing), and expressive aesthetics referring to creativity and originality Hartmann_De. Mahlke and Thüring Mahlke_Thüring_2007 argue that classical aesthetics is perceived more evenly, whereas expressive aesthetics varies with framing; as such, in this paper, we focus on classical aesthetics. For a more complete review of the literature, refer to Ahmed et al. Ahmed_Al.
The relationship between UI aesthetics and functionality has been investigated by a number of authors. Some argue that the two goals often conflict Karvonen_2000, whereas Kurosu and Kashimura kurosu1995apparent found a positive correlation between the beautifulness of ATM interfaces and their usability. Due to the subjective nature of aesthetics, the strength of the correlation coefficient has been shown to be culturally dependent tractinsky1997aesthetics. To our knowledge, no previous publication has investigated aesthetics in the context of CVD. While our simulation-based experimental methodology is not robust enough to measure aesthetics for CVD people, it allows us to hypothesize that functionality and classical aesthetics correlate in CVD-simulated UIs, and that classical aesthetics could be negatively impacted by automatic invasive techniques such as the high contrast mode.
Figure 3: Illustration of a website automatically adapting to Windows high contrast mode (Night Sky). The reduced color palette is designed to increase contrast between key UI components.
3 Materials and Methods
In our experiment, we aim to comparatively study the following:
  • Functionality: how well could a CVD user use the interface? I.e., compared to a non-CVD user, is usability reduced? Are key visual elements equally perceivable?
  • Aesthetics: according to classical aesthetics, would the clarity and orderliness of the UI be retained when viewed by a CVD observer?
As discussed earlier, the experimental design is inherently non-trivial. Individuals with CVD cannot make a comparative judgment on whether the stimulus they perceive has the same functionality and aesthetics as a stimulus that they could never see. Conversely, an individual with no CVD does not have sufficient experience to rate how an individual with CVD might perceive a stimulus. A large-scale study with cross-population comparisons of subjective ratings might be considered; however, the subjective nature of aesthetics and the various experiences and consequent biases of the two populations might still prevent such studies from drawing reliable conclusions.
Therefore, we instead build on the successful field of physiologically-based models of CVD simulation that allow non-CVD observers to inspect UIs as seen by CVD observers. Such models are based on a strong scientific understanding of early vision and have been verified by controlled user experiments Machado2009. An inherent limitation of this approach is non-CVD observers’ bias towards the more familiar reference images; therefore, we do not consider the protocol to be suitable to judge whether the functionality and aesthetics would be fully retained for a CVD observer. However, we argue that relative scores across applications and CVD mitigation strategies still reveal crucial insights. Furthermore, a simulation-based pipeline has the benefit that it is easily deployable in a commercial environment.
For an overview of the methodology, see Figure LABEL:fig:methods-fc. In the following subsections, we describe the stimuli, setup, task, participant sample set, and the statistical methods applied.